# Generative Consistency Models for Estimation of Kinetic Parametric Image Posteriors in Total-Body PET

### v1

@@ 2.3.3 Performance Metrics
Method performance was quantified with three statistics. First, the dissimilarity between the posterior produced by method $m$, $q_{m}(\theta|y)$, and the reference MCMC posterior, $p_{MCMC}(\theta|y)$, was measured by an empirical Kullback–Leibler divergence,
@@ 2.4 Dynamic [18F]FDG Total-Body Study
List-mode data were histogrammed into 35 dynamic frames, each with matrix size $220\times 220\times 645$. Normalization, random, attenuation, decay and scatter corrections were applied and the data were reconstructed using ordinary Poisson ordered subsets expectation maximization algorithm (OP-OSEM, 8 iterations, 5 subsets), with time-of-flight weighting and a maximum ring difference of 322. Spatiotemporal HYPR-LR filtering [31] was applied to the dynamic series, reducing noise by pooling information across space and time while preserving high-frequency spatial detail and temporal dynamics for subsequent kinetic analysis. The AIF was extracted from the ascending aorta. First, a three-dimensional mask of the aorta was produced on the native CT images using MOOSE segmentation [32], and the mask was then morphologically eroded (two-voxel radius) to minimize partial volume contamination. The whole-blood AIF was taken as the mean activity concentration within this eroded region for each frame. Using the AIF and reconstructed voxel-based TACs as input, posteriors of the micro-parameters, $K_{1},k_{2},k_{3}$ and $k_{4}$ were estimated using CM. Then, the net influx rate, $K_{i}$, was computed for every voxel as $K_{i}=K_{1}k_{3}/(k_{2}+k_{3})$ to generate parametric images of $K_{i}$ which were compared with voxel-based Patlak $K_{i}$ images.
@@ 3.1 Evaluation on Simulation Data
We first evaluated our proposed CM method on testing data. Fig. 3 shows an example where the posteriors were estimated on one TAC-AIF pair from the testing dataset. The CM yields posteriors that almost overlap with those generated by the MCMC reference method for kinetic parameters except $k_{3}$. Its modes coincide and its spread appears within $5\%$ of the MCMC widths, suggesting minimal bias and good variance fidelity. For $k_{3}$, the CM posterior is left-shifted and slightly under-dispersed relative to the MCMC reference, and the competing methods display a comparable tendency. The ABC posterior retains similar means but displays noticeably larger tails, indicating larger variance. Among the diffusion based methods, the DDPM produces slightly wider peaks, particularly for $K_{1}$, $k_{2}$, and $k_{4}$, whereas the SBD exhibits narrower shoulders, suggesting slight overdispersion and underdispersion, respectively. The CVAE compresses each distribution, implying a tendency to underestimate variance for $K_{1}$, $k_{2}$, and $k_{4}$, while the GAN shifts the means of all kinetic parameters and narrows their peaks, suggesting both bias and decreased uncertainty.
@@
Fig. 4 extends the analysis of Fig. 3 to 100 independent TAC–AIF pairs. For every pair we compared the posteriors produced by our CM with those from MCMC and five competing methods—ABC, DDPM, SBD, CVAE and GAN. Panel (a) plots the Kullback–Leibler divergence (KLD) between each method’s posterior and the MCMC reference. The CM attained the lowest KLD median for $K_{1}$, the second lowest for $k_{2}$ and $k_{3}$, and yielded a KLD for $k_{4}$ comparable to ABC, outperforming both the CVAE and GAN. Panel (b) focuses on point-estimate accuracy using the absolute percentage error (APE) between the posterior mean and the ground-truth parameters. We include MCMC in this comparison as a benchmark. Because each posterior is conditioned on one noisy TAC–AIF realization, even MCMC yields a non-zero APE—the posterior mean need not coincide exactly with the latent kinetic vector. Consequently, boxplots that cluster near the MCMC box denote methods whose bias is comparable to MCMC. In this setting, the error distribution of CM most closely approximates that of the reference MCMC method for $K_{1}$, $k_{2}$ and $k_{3}$, indicating minimal bias in the posterior mean estimates. However, for $k_{4}$, no approach achieves errors near the MCMC benchmark; instead, diffusion‐based methods (DDPM and SBD) outperform CM. Panel (c) examines uncertainty estimation, reporting for each kinetic parameter the percentage error between a method’s posterior standard deviation and that obtained with MCMC. Accurate uncertainty estimation is critical, as downstream model selection depends on reliable credible intervals. For $K_{1}$, the CM and the diffusion‐based methods (DDPM, SBD) exhibit median absolute errors of approximately $20\%$, whereas all other approaches show larger medians. A similar pattern holds for $k_{2}$: DDPM and SBD remain closest to MCMC, with CM displaying a slightly higher median error. In the case of $k_{3}$, ABC achieves the lowest median error, and DDPM and SBD follow within less than one tenth, while CVAE and GAN overshoot by roughly twice that margin. Uncertainty estimation proves most demanding for $k_{4}$. SBD yields the best alignment with MCMC, followed by CM. DDPM and ABC produce comparable median errors, and CVAE and GAN deviate the most.
Figure 5: SUV reference image and $K_{i}$ parametric maps from a dynamic TB-PET data. From left to right: static clinical SUV image; conventional Patlak $K_{i}$ map; $K_{i}$ map estimated by the proposed consistency model (CM); $K_{i}$ uncertainty map (i.e., width of 95% credible interval). Six representative voxels are marked on the SUV image by green dots in the left frontal cortex, myocardium, tumour, stomach, left kidney, and tensor fasciae latae. Posterior $K_{i}$ distributions for these voxels are shown in the inset density plots.
@@ 3.2 Total-Body Parametric Imaging
Fig. 5 presents the SUV and $K_{i}$ images reconstructed with CM and Patlak analysis. The SUV image shows the expected FDG biodistribution: high uptake of FDG in cortical brain and myocardium, and renal cortices, and excretion into the urinary bladder. Intermediate uptake is observed throughout the liver and skeletal muscle; and very low uptake in the lungs. A focal area of high uptake is evident at the inferior right hepatic margin corresponding to a known metastatic tumour. Qualitatively, the Patlak map preserves the expected physiological pattern of high $K_{i}$ in cortical grey matter of the brain, myocardium and renal cortex; intermediate uptake in liver and skeletal muscle, and low activity in lung parenchyma, but exhibits pronounced streak artefacts and amplified noise in low count regions.
@@
The CM derived $K_{i}$ image (third panel of Fig. 5) is in good agreement with the Patlak image. High influx rates are clearly delineated in the same organs and tissues as the Patlak $K_{i}$ image. The hyper‑metabolic tumour in the right lobe of the liver is evident but without the streak artefacts and amplified noise seen in the Patlak image. Across low‑count regions like distal extremities, pelvis and soft tissues, the CM image exhibits reduced variance, indicating a more favourable noise–bias trade‑off than Patlak. The accompanying $K_{i}$ uncertainty map (Fig. 5 fourth panel) depicts, for every voxel, the width of the 95% credible interval derived from the voxel-wise posterior distributions, elucidating the spatial distribution of estimation uncertainty. The highest uncertainty is confined to regions exhibiting intense tracer activity and rapid kinetics, principally the renal cortices and urinary bladder where the 2-tissue compartment model does not apply. In contrast, the cerebral cortex and skeletal muscles demonstrate low uncertainty, and the tumour in the liver likewise shows low levels of uncertainty, indicating a high degree of confidence in the parameter estimates for these regions. Such explicit, spatially resolved confidence information is unavailable from SUV or Patlak analyses and constitutes a key advantage of the proposed CM framework.
@@
CM derived $K_{i}$ posteriors from six representative voxels are selected and compared to posteriors predicted by MCMC and ABC, as shown in Fig. 5 fifth panel. For voxels from cerebral cortex, myocardium and the liver tumour, CM posteriors almost coincide with the MCMC reference, whereas ABC exhibits broader tails, suggesting an overestimation of uncertainty. In the stomach voxel CM matches MCMC, while the ABC posterior is right skewed and its longer right tail yields a smaller posterior mean. In the right kidney, ABC and CM align with MCMC. In the tensor fasciae latae (a muscle that attaches to the iliac crest of the pelvis), the CM posterior peaks at a slightly higher $K_{i}$ value than the MCMC posterior exhibiting a comparable but marginally broader spread, while the ABC posterior shares the MCMC peak location yet displays a markedly elongated right tail.
@@ 3.3 Voxel-wise Model Selection
To complement the $K_{i}$ maps we used the posterior samples produced by the CM to create a voxel-wise binary parametric image of kinetic reversibility based on the efflux rate $k_{4}$. To generate this image, we assigned a 0 or 1 to each voxel indicating conformance with the irreversible ($k_{4}=0$) and reversible ($k_{4}>0$) two compartment models, respectively, based on the posterior probability that the efflux rate $k_{4}$ for the corresponding voxel is statistically distinguishable from zero. Specifically, for each voxel we examined the 95% credible interval (CrI) of the posterior for $k_{4}$. If the interval included zero or if zero was on the right side of the CrI, the voxel was labelled irreversible. If the entire interval lay above zero, we classified the voxel as reversible.
Clinically, this map helps distinguish tissues in which tracer becomes metabolically trapped, such as cortical brain, myocardium and malignant tumours, from those where tracer exchanges reversibly with the free tissue compartment. Unlike point-estimation methods such as Patlak or nonlinear least squares, CM outputs a full posterior that can support model selection.
@@
Figure 6: Voxel-wise model selection map for the efflux rate $k_{4}$. A binary mask on the CT volume labels voxels with a high posterior probability of irreversible efflux ($k_{4}=0$; black) while leaving reversible voxels transparent. Posterior $k_{4}$ distributions from CM (blue) and MCMC (red) are shown for six example voxels.
@@
In Fig. 6 (left side), brain, myocardium and skeletal muscle (indicated by the tensor fasciae latae) appear to be predominantly irreversible. On the contrary, healthy liver and renal cortex are classified as reversible. In particular, most liver voxels are classified as reversible, while the tumour in the inferior right lobe (previously highlighted in the $K_{i}$ map) shows a contiguous cluster of irreversible voxels. Posteriors for six representative voxels illustrate the classification process (Fig. 6 right side). In cerebral cortex and myocardium the posteriors of $k_{4}$ predicted by CM are centred on the positive side of zero but the $95\%$ CrI contained the origin, hence both voxels are labelled as irreversible. The tumour voxel in the inferior right liver shows a similar slightly left-shifted probability density whose interval still contains zero, and is classified as irreversible. By contrast, the voxel in the stomach and in the right kidney have posteriors that sit on the positive side of the axis, placing zero outside their $95\%$ CrIs, thus classifying them as reversible. Finally, the tensor fasciae latae voxel exhibits a $95\%$ CrI containing zero, leading to an irreversible classification. Across all six voxels, the CM posteriors broadly mirror the distributions obtained by MCMC.

### v2

@@ 2.3.3 Performance Metrics
First, the dissimilarity between the posterior produced by method $m$, $q_{m}(\theta|y)$, and the reference MCMC posterior, $p_{MCMC}(\theta|y)$, was measured by an empirical Kullback–Leibler divergence,
@@ 2.4 Dynamic [18F]FDG Total-Body Study
List-mode data were histogrammed into 35 dynamic frames, each with matrix size $220\times 220\times 645$. Normalization, random, attenuation, decay and scatter corrections were applied and the data were reconstructed using ordinary Poisson ordered subsets expectation maximization algorithm (OP-OSEM, 8 iterations, 5 subsets), with time-of-flight weighting and a maximum ring difference of 322. Spatiotemporal HYPR-LR filtering [35] was applied to the dynamic series. The AIF was extracted from the ascending aorta. First, a three-dimensional mask of the aorta was produced on the native CT images using MOOSE segmentation [36], and the mask was then morphologically eroded (two-voxel radius) to minimize partial volume contamination. The whole-blood AIF was taken as the mean activity concentration within this eroded region for each frame. Previous validation studies have demonstrated that ascending aorta IDIFs obtained with erosion strategies provide robust kinetic estimates effectively equivalent to arterial sampling [37]. Using the AIF and reconstructed voxel-based TACs as input, posteriors of the micro-parameters, $K_{1},k_{2},k_{3}$ and $k_{4}$ were estimated using CM. Then, the net influx rate, $K_{i}$, was computed for every voxel as $K_{i}=K_{1}k_{3}/(k_{2}+k_{3})$ to generate parametric images of $K_{i}$ which were compared with voxel-based Patlak $K_{i}$ images.
@@ 3.1 Evaluation on Simulation Data
We first evaluated our proposed CM on simulated testing data. Fig. 3 shows an example where the posteriors were estimated on one TAC-AIF pair from the testing dataset. The CM yields posteriors that closely approximate the ones predicted by the MCMC, particularly for $K_{1}$. For the other parameters ($k_{2},k_{3},k_{4}$), while minor distributional shifts or variances are observed compared to the MCMC prediction, the CM posterior consistently encapsulates the ground truth values (dashed lines) within its credible interval (CrI). The ABC posterior retains similar means but displays noticeably larger tails, indicating larger variance. Among the diffusion based methods, the DDPM produces slightly wider peaks, particularly for $K_{1}$, $k_{2}$, and $k_{4}$, whereas the SBD exhibits narrower shoulders, suggesting slight overdispersion and underdispersion, respectively. The CVAE compresses each distribution, implying a tendency to underestimate variance for $K_{1}$, $k_{2}$, and $k_{4}$, while the GAN shifts the means of all kinetic parameters and narrows their peaks, suggesting both bias and decreased uncertainty.
@@
Fig. 4 extends the analysis of Fig. 3 to 100 independent TAC–AIF pairs. For every pair we compared the posteriors produced by our CM with those from MCMC and five baselines—ABC, DDPM, SBD, CVAE and GAN. Panel (a) plots the Kullback–Leibler divergence (KLD) between each method’s posterior and the MCMC reference. The CM attained the lowest KLD median for $K_{1}$, the second lowest for $k_{2}$ and $k_{3}$, and yielded a KLD for $k_{4}$ comparable to ABC, outperforming both the CVAE and GAN. Panel (b) focuses on point-estimate accuracy using the absolute percentage error (APE) between the posterior mean and the ground-truth parameters. We include MCMC in this comparison as a benchmark. Because each posterior is conditioned on one noisy TAC–AIF realization, even MCMC yields a non-zero APE—the posterior mean need not coincide exactly with the latent kinetic vector. Consequently, boxplots that cluster near the MCMC box denote methods whose bias is comparable to MCMC. In this setting, the error distribution of CM most closely approximates that of the reference MCMC method for $K_{1}$, $k_{2}$ and $k_{3}$, indicating minimal bias in the posterior mean estimates. However, for $k_{4}$, no approach achieves errors near the MCMC benchmark; instead, diffusion‐based methods (DDPM and SBD) outperform CM. Panel (c) examines uncertainty estimation, reporting for each kinetic parameter the percentage error between a method’s posterior standard deviation and that obtained with MCMC. Accurate uncertainty estimation is critical, as downstream model selection depends on reliable CrIs. For $K_{1}$, the CM and the diffusion‐based methods (DDPM, SBD) exhibit median absolute errors of approximately $20\%$, whereas all other approaches show larger medians. A similar pattern holds for $k_{2}$: DDPM and SBD remain closest to MCMC, with CM displaying a slightly higher median error. In the case of $k_{3}$, ABC achieves the lowest median error, and DDPM and SBD follow within less than one tenth, while CVAE and GAN overshoot by roughly twice that margin. Uncertainty estimation proves most demanding for $k_{4}$. SBD yields the best alignment with MCMC, followed by CM. DDPM and ABC produce comparable median errors, and CVAE and GAN deviate the most.
To assess the robustness of the CM to variable tracer arrival times, we stratified the estimation error for kinetic parameters ($K_{1},k_{2},k_{3},k_{4}$) against the ground-truth delay $t_{0}$ present in the test samples. As shown in Fig. 5, the APE remains stable across the delay range (0 to $1.25$ min) for all parameters, confirming that the CM successfully marginalizes over the delay parameter without requiring explicit temporal alignment of the AIF.
Figure 5: Robustness of Kinetic Parameter Estimation to Tracer Arrival Delays. Analysis of estimation accuracy with respect to tracer arrival delay ($t_{0}$). The absolute percentage error (APE) for each kinetic parameter is plotted against the ground-truth delay time for test samples.
3.2 Evaluation on Clinical Data
The proposed CM was assessed on the clinical [18F]FDG total-body PET scan described in Section 2.4. To demonstrate clinical relevance across a range of organs, we evaluated posterior inference in four distinct targets: cerebral gray matter, myocardium, a hepatic tumour, and healthy liver. As illustrated in Fig. 6 rows 1–4, the CM posterior distributions (blue) exhibit agreement with the MCMC reference (red) across all kinetic parameters. In high-signal regions such as the cerebral cortex and myocardium, the model consistently resolves the sharp posterior modes associated with rapid tracer delivery and uptake. In the hepatic tumour, the CM accurately captures the irreversible trapping characteristic of malignant lesions, showing high probability density for $k_{4}=0$. For the healthy liver background, the CM correctly differentiates the distinct metabolic kinetics (lower $k_{3}$, higher $k_{4}$) from the adjacent tumour tissue. To substantiate these observations quantitatively, Fig. 6 row 5 presents the KLD between posteriors predicted by CM and MCMC, computed with 100 voxels for each region. The results indicate consistently low divergence (median KLD generally <0.5<0.5) across distinct physiological environments, from the high-noise tumour to the high-uptake myocardium. This confirms that the CM’s predictive accuracy is not an artifact of selected examples but reflects a robust generalized ability to recover the posterior distribution.
Figure 6: Evaluation of voxel-wise posterior inference in clinical tissues. Rows 1–4: Posterior distributions of kinetic parameters ($K_{1},k_{2},k_{3},k_{4}$) for representative single voxels extracted from four regions: Brain (cerebral cortex), Myocardium, Tumour (hepatic lesion), and healthy Liver. Row 5: Boxplots of the KLD between posteriors predicted by CM and MCMC, calculated over 100 voxels within each region.
@@ 3.3 Total-Body Parametric Imaging
Fig. 7 presents the SUV and $K_{i}$ images calculated with Patlak analysis, nonlinear least squares (NLS), and the CM. The SUV image shows the expected FDG biodistribution: high uptake of FDG in cortical brain and myocardium, and renal cortices, and excretion into the urinary bladder. Intermediate uptake is observed throughout the liver and skeletal muscle; and very low uptake in the lungs. A focal area of high uptake is evident at the inferior right hepatic margin corresponding to a known metastatic tumour. Qualitatively, the Patlak map preserves the expected physiological pattern of high $K_{i}$ in cortical grey matter of the brain, myocardium and renal cortex; intermediate uptake in liver and skeletal muscle, and low activity in lung parenchyma, but exhibits pronounced streak artefacts and amplified noise in low count regions.
Figure 7: SUV reference image and $K_{i}$ parametric maps from a dynamic TB-PET data. From left to right: static clinical SUV image (average of the last frame, 45–50 min); conventional Patlak $K_{i}$ map (linear fit over $t^{*}>18$ min, last 8 frames); $K_{i}$ map calculated by fitting the full two tissue compartment models using nonlinear least squares; $K_{i}$ map estimated by the proposed consistency model (CM); $K_{i}$ uncertainty map (i.e., width of 95% CrI).
@@
The CM derived $K_{i}$ image generally agrees with the Patlak image. High influx rates are clearly delineated in the same organs and tissues as the Patlak $K_{i}$ image. The hyper‑metabolic tumour in the right lobe of the liver is evident but without the streak artefacts and amplified noise seen in the Patlak image. Across low‑count regions like pelvis and soft tissues, the CM image exhibits reduced variance, indicating a more favorable noise–bias trade‑off than Patlak. To investigate the differences between CM and Patlak, we computed ROI-averaged TACs and inferred the kinetic parameters using the Patlak, CM and MCMC. Patlak yields lower estimates compared to the MCMC reference. In the brain ($K_{i}^{\text{Patlak}}{=}0.0233$, $K_{i}^{\text{MCMC}}{=}0.0347$) and myocardium ($K_{i}^{\text{Patlak}}{=}0.0733$, $K_{i}^{\text{MCMC}}{=}0.0912$), the Patlak estimates are notably lower, whereas the CM ($0.0369$ and $0.0881$, respectively) is closer to the MCMC reference. This distinction is also critical in the liver where the Patlak predicts $K_{i}=0.0040$ while the CM ($0.0090$) is closer to the truth predicted by MCMC ($0.0078$). This highlights Patlak’s limitations: assuming irreversible uptake ($k_{4}=0$) while neglecting fractional blood volume ($v_{b}$) and tracer arrival delay, thereby leading to biased estimates. For comparison with a standard full-model approach, voxel-wise NLS estimation of the reversible two tissue compartment model was also performed. While NLS aligns generally with the CM estimates in high-count regions (e.g., brain, heart, kidneys), it exhibits higher variance in lower-count regions (e.g., liver, muscle) compared to the CM posterior means. The accompanying $K_{i}$ uncertainty map depicts, for every voxel, the width of the 95% CrI derived from the voxel-wise posterior distributions, elucidating the spatial distribution of estimation uncertainty. The highest uncertainty is confined to regions exhibiting intense tracer activity and rapid kinetics, principally the renal cortices and urinary bladder where the 2-tissue compartment model does not apply. In contrast, the cerebral cortex and skeletal muscles demonstrate low uncertainty, and the tumour in the liver likewise shows low levels of uncertainty, indicating a high degree of confidence in the parameter estimates for these regions. Such explicit, spatially resolved confidence information is unavailable from SUV, NLS or Patlak analyses and constitutes a key advantage of the proposed CM framework.
@@
CM derived $K_{i}$ posteriors from six representative voxels are selected and compared to posteriors predicted by MCMC and ABC, as shown in Fig. 8. For voxels from cerebral cortex, myocardium and the liver tumor, CM posteriors almost coincide with the MCMC reference, whereas ABC exhibits broader tails, suggesting an overestimation of uncertainty. In the stomach voxel CM matches MCMC, while the ABC posterior is right skewed and its longer right tail yields a smaller posterior mean. In the right kidney, ABC and CM align with MCMC. In the tensor fasciae latae (a muscle that attaches to the iliac crest of the pelvis), the CM posterior peaks at a slightly higher $K_{i}$ value than the MCMC posterior exhibiting a comparable but marginally broader spread, while the ABC posterior shares the MCMC peak location yet displays a markedly elongated right tail.
Figure 8: Posterior predictions of $K_{i}$ in selected regions. Six representative voxels are marked on the SUV image in Fig. 7 by green dots in the left frontal cortex, myocardium, tumor, stomach, left kidney, and tensor fasciae latae. Posterior $K_{i}$ distributions predicted by MCMC, ABC and CM for these voxels are shown in the density plots.
Figure 9: Voxel-wise estimation of kinetic micro-parameters. Total-body posterior mean maps for $K_{1}$ (delivery) and $k_{4}$ (clearance) generated by the CM.
Fig. 9 shows voxel-wise posterior mean maps for micro-kinetic parameters $K_{1}$ and $k_{4}$. The estimated values align with the established physiological kinetics. The $K_{1}$ map specifically shows high-perfusion regions, exhibiting maximal intensity in the kidneys, myocardium, and liver. In the $k_{4}$ map, the model distinguishes metabolic phenotypes. The hepatic tumour, myocardium, and cerebral gray matter exhibit negligible clearance ($k_{4}\approx 0$) consistent with high FDG trapping, whereas the healthy liver displays non-zero dephosphorylation rates.
@@ 3.4 Voxel-wise Model Selection
We used the posterior samples produced by the CM to create a voxel-wise binary parametric image of kinetic reversibility based on the efflux rate $k_{4}$. To generate this image, we assigned a 0 or 1 to each voxel indicating conformance with the irreversible ($k_{4}=0$) and reversible ($k_{4}>0$) two compartment models, respectively, based on the posterior probability that the efflux rate $k_{4}$ for the corresponding voxel is statistically distinguishable from zero. Specifically, for each voxel we examined the 95% CrI of the posterior for $k_{4}$. If the interval included zero or if zero was on the right side of the CrI, the voxel was labelled irreversible. If the entire interval lay above zero, we classified the voxel as reversible. Clinically, this map helps distinguish tissues in which tracer becomes metabolically trapped, such as cortical brain, myocardium and malignant tumours, from those where tracer exchanges reversibly with the free tissue compartment. Unlike point-estimation methods such as Patlak or nonlinear least squares, CM outputs a full posterior that can support model selection.
@@
Figure 10: Comparison of voxel-wise model selection. A binary mask labels voxels with a high posterior probability of irreversible efflux ($k_{4}=0$ black, $k_{4}\neq 0$ white). Left: CM-derived model selection map (based on posterior $k_{4}$ support). Middle: Posterior $k_{4}$ distributions from CM (blue) and MCMC (red) are shown for six example voxels. Right: NLS-derived model selection map (based on BIC).
@@
In Fig. 10 (left panel), brain, myocardium and skeletal muscle (indicated by the tensor fasciae latae) appear to be predominantly irreversible. On the contrary, healthy liver and renal cortex are classified as reversible. In particular, most liver voxels are classified as reversible, while the tumour in the inferior right lobe (previously highlighted in the $K_{i}$ map) shows a contiguous cluster of irreversible voxels. Posteriors for six representative voxels illustrate the classification process (Fig. 10 middle panel). In cerebral cortex and myocardium the posteriors of $k_{4}$ predicted by CM are centred on the positive side of zero but the $95\%$ CrI contained the origin, hence both voxels are labelled as irreversible. The tumour voxel in the inferior right liver shows a similar slightly left-shifted probability density whose interval still contains zero, and is classified as irreversible. By contrast, the voxel in the stomach and in the right kidney have posteriors that sit on the positive side of the axis, placing zero outside their $95\%$ CrIs, thus classifying them as reversible. Finally, the tensor fasciae latae voxel exhibits a $95\%$ CrI containing zero, leading to an irreversible classification. Across all six voxels, the CM posteriors broadly mirror the distributions obtained by MCMC.
We compared our posterior-based model selection against standard frequentist approach. We calculated the Bayesian Information Criterion (BIC) for competing reversible and irreversible models using voxel-wise NLS estimator. As shown in Fig. 10 right panel, while BIC yields similar classifications in high-count organs (e.g., myocardium), it demonstrates instability in heterogeneous lesions. Specifically, NLS-BIC and CM showed inconsistencies in liver tumors. CM identified irreversible uptake kinetics in tumors ($k_{4}\approx 0$), and according to literature studies, the uptake of FDG by these tumors is more likely to be irreversible [38, 39, 40]. This underscores that direct inspection of the posterior CrI is not only more computationally efficient (avoiding multi-model fitting) but also reliable in low-SNR regions of voxel-wise TB-PET.